One number is 3 less than another. If 4 times the smaller number minus 3 times the larger number is 4, find the two numbers.
Answer by mananth(16946)

let one number be x
the other number = x-3.
..
4(x-3)-3x = 4
4x-12-3x=4
x=16
the other number is 13